calculate the molarities of the solutions prepared by dissolving the following solutes in sufficient water to produce the given volumes of solution.
16.7 g CaCl^2 in 400.0 mL solution
1
Expert's answer
2017-05-22T08:18:20-0400
Answer:
n(CaCl2) = m(CaCl2) / M(CaCl2) = 16.7 / 111 = 0.15
c(CaCl¬2) = n(CaCl2) / V(solution) = 0.15 / 0.4 = 0.375 M